(1) a) Give parametric equation (point-direction form) of the line which lies on both of the planes: x+y+z= 1 and -x+2y + 10z 2. What is the direction d of this line? b) Let ny and n₂ be the normal vectors to the two given planes. Without actual computation, describe the relationship between d and n₁ x n₂.


1
Expert's answer
2021-09-26T18:11:35-0400

a) "n_1=\langle 1, 1, 1\rangle, n_2=\langle -1, 2, 10\rangle"


"n_1\times n_2=\begin{vmatrix}
 i & j & k \\
 1 & 1 & 1 \\
 -1 & 2 & 10 \\
\end{vmatrix}"

"=i\begin{vmatrix}
 1 & 1 \\
 2 & 10
\end{vmatrix}-j\begin{vmatrix}
 1 & 1 \\
 -1 & 10
\end{vmatrix}+k\begin{vmatrix}
 1 & 1 \\
 -1 & 2
\end{vmatrix}"

"=8i-11j+3k"

"d=\langle 8, -11, 3\rangle"

"x+y+z= 1""-x+2y + 10z =2"

"x+y+z= 1""3y+11z= 3"

"x_1=0, y_1=1, z_1=0"


"x_2=8, y_2=-10, z_2=3"

"d=\langle x_2-x_1, y_2-y_1, z_2-z_1\rangle"

"d=\langle 8-0, -10-1, 3-0\rangle"

"d=\langle 8, -11, 3\rangle"

b)


"d=n_1\times n_2"

"d\perp n_1, d\perp n_2"